# Studying Differential Mental Health Expressions in India

### v2

@@ 4.1.2 LIWC
Fifty-two LIWC categories were significantly associated ($p<0.05$) with the MH-India group, whereas sixty categories were found to be correlated with the MH-ROW group. We provide the Top 20 LIWC categories for both groups in Table 2. While negative tone, anxiety, and personal pronouns are correlated with depression in both groups, sadness, an additional affect is uniquely seen in the MH-Indian group. Depression-related discourse in MH-ROW group is past-focused, which is a known marker of depression whereas it is present-focused in India. Social behavioral attributes (communication, politness), work, and achievement are correlated with depression in the MH-India group but not in MH-ROW. Somatic symptoms/illness (pain, sick) are seen in the MH-India group, whereas substance abuse/addiction is correlated with the MH-ROW group.
MH-India MH-ROW
Theme Top Words r Theme Top Words r
Negative/ Suicidal thoughts life, parents, family, hate, die 0.280 Negative Feelings feel, myself, feeling, depression, anymore 0.358
Depression mentions anymore, depression, tired, depressed, everyday 0.229 Negative Feelings went, didn, crying, mad, stayed 0.340
Negative Feelings feel, myself, feeling, depression, anymore 0.227 Depression mentions anymore, depression, tired, depressed, everyday 0.338
Friends & Relationships friends, talk, social, anxiety, alone 0.226 Negative Feelings feeling, body, feels, heart, scared 0.322
Friends & Relationships love, heart, loved, beautiful, happiness 0.207 Symptoms sick, woke, stomach, switched, asleep 0.318
Table 3: Top Words and Correlation Coefficient (r) by Theme for MH-India and MH-ROW
@@ 4.1.3 LDA Topics
Of $2000$ topics, $109$ were found to be significant ($p<0.05$) for the MH-India group and $216$ for MH-ROW group. The top topics and the corresponding Pearson r, are provided in Table LABEL:tab:topics_output. The most prevalent topics in the MH-India group discussed negative/suicidal thoughts (life, parents, family, hate, die), academic and job stressors ((college, exam, study, university, engineering), job, degree, college, school, career) and family and relationships (love, heart, loved, beautiful, happiness). In contrast, the prevalent themes in MH-ROW are negative feelings and emotions ((feel, myself, feeling, depression, anymore);(went, didn, crying, mad, stayed); (feeling, body, feels, heart, scared)). Topics such as mental disorders (anxiety, depression, mental, medication, disorder), goodbyes (sister, ma, papa, clutching, plane) and anger (hate, angry, rant, respect, ugly) are common in both sets.
@@ 4.2 Validation
To what extent the open vocabulary topics having a significant correlation with the MH-India group are prevalent in Indian patients? - A Likert scale of 0-5 is provided where $5$ indicates ‘Highly Prevalant’ and ‘0’ indicates ‘Not observed at all’.
@@ 4.3 Is the language in MH-India different from control groups?
We used logistic regression model with each feature set (i.e., n-grams, LIWC and LDA topics) to predict membership in MH-India vs MH-ROW. High AuC scores (See Table 4) demonstrate that users’ language in the MH-India group significantly differs from those in the control groups, including MH-ROW. All language feature groups (i.e., n-grams, LIWC, and LDA topics) have fairly high AUC, indicating differences in depression expression at 1-3 gm level, in psycho-social categories as well as in latent topics of discussion.
@@ 5 Discussion
Our work reveals significant differences in the language markers of mental health expressions of Indian users compared to those from outside India. The association with politeness coupled with discussions around family, work and achievement indicates that the users in the MH-India group tend to associate mental health with their ability and social relationships as opposed to more mentions of swear words and feelings in the ROW group. Academic and family pressures are unique to Indian users, possibly due to the collectivist nature of Indian society Chadda and Deb (2013). The matching performed across control groups rules out the majority young demographic in social media data as the possible reason. The association with present focus words in MH-India contrasts with the widespread belief of self-focussed rumination when suffering from mental illness Park et al. (2017). MH-India group also tends to reason (‘causation’ words - how, because, why) more in their language in contrast to expressing feelings in MH-ROW group. Communication (phone, call, message, post/tweet/meme, sms/texting, chat) is also exclusive to the MH-India group. We speculate high reliance on social media platforms for mental health support and privacy potentially due to stigma associated with mental health in face-to-face conversations Shidhaye and Kermode (2013).
Our work reaffirms cultural differences such as substance use in the MH-ROW population vs somatic complaints in the MH-India population. Unlike in Western cultures where substance use is a barrier to mental health support, it is familial concerns and the fear of stigma in MH-India Biswas et al. (2016). We speculate based on these findings that culturally-aware mental health intervention techniques should seek to bridge the gap to treatment by addressing hypochondriacal ideas and familial embarrassment in particular.
@@
Only $56\%$ of 109 topics correlated with the MH-India group were labeled as prevalent in Indian patients by clinical psychologists. We speculate that some of the topics not labeled as "prevalent" are unseen or emerging themes. Of Top-20 topics in MH-India, "not prevalent" topics revolve around Video Games/Online Content, Grooming/Physical Appearance, and Programming, indicating the influence of digital content, growing isolation, and low self-esteem amongst the undiagnosed young population. These topics could be underrecognized concerns. The second group of topics includes Environmental Impact of Energy Sources, Humorous content and reactions, among others. Previous research has suggested that people (particularly young people) are increasingly climate anxious and that humor on social media is often used to cope with mental health challenges Schneider (2018); Sanson (2022). Furthermore, certain pop culture references may be crucial to understanding the narrative - as one clinical psychologist pointed out, the top word "Singh" in one of the topics may correspond to the suicide of late Bollywood star Sushant Singh Rajput, a significant event that potentially catalyzed a range of important conversations surrounding mental health in India Akbar et al. (2020).
@@
The growing treatment gap for mental disorders is a major concern in Indian society. The economic loss from mental health conditions between 2012-2030 is estimated at USD 1.03 trillion [^2] United Nations: https://www.who.int/india/health-topics/mental-health. Automated systems that could diagnose and support mental well-being can potentially alleviate the lack of resources, but they would only be useful when designed considering the cultural sensitivities and norms of society. This study establishes significant linguistic variations in the mental health-related language in social media posts by Indians compared to individuals from the rest of the world. It thus becomes imperative to test the computational depression detection models for cultural invariance before their deployment in clinical settings Aguirre et al. (2021); Rai et al. (2023) and indicate the need for socio-culturally aware mental health models to prevent misdiagnosis.
@@ 6 Background
Depression and anxiety disorders are the most imminent mental health challenges, with the highest contribution to Indian Disability Adjusted Life Years  Sagar et al. (2020). Fear of embarrassing family members is a primary barrier to mental health recovery in India whereas, for example, substance abuse is the major hurdle in America  Biswas et al. (2016). $71\%$ of Indians exhibited stigma when answering questions about mental health Foundation (2018). Relatedly, somatic symptoms, hypochondriasis, anxiety, and agitation are more commonly seen in Indian patients compared to psychological symptoms  Gada (1982). While the extent of mental health stigma and treatment (un)availability is often studied, it remains unknown how individuals suffering from mental disorders in India express and seek support on social media.
There is accumulating evidence that suggests language markers of depression vary with demographics such as race Rai et al. (2024); Aguirre and Dredze (2021), immigrant status Mittal et al. (2023a), and geographic location De Choudhury et al. (2017). For instance, Indian and South Africa-based users are less candid in their posts and less likely to exhibit negative emotions in comparison to their Western counterparts De Choudhury et al. (2017). Another study studied the language of Asian users including Indian, Malaysian, and Filipino on Mental Health Support Forum, Talklife Pendse et al. (2019) and found that Indians discuss “wanting or needing friends” more than other countries. .
@@
Our university’s Institutional Review Board deemed this study exempt due to the public nature of all data. While Reddit data is public, it may contain users’ personal information, including city and town. We limited our analysis to country and state-level geolocation information to reduce the possibility of personally identifying individuals. Gender was predicted using a continuous scale, with extremes indicating masculinity and feminity. We exercised caution while presenting linguistic patterns and examples not to reveal any individual’s timeline quotes. Members of our team have not viewed or worked with individual-level granular data. When done ethically with respect to user anonymity and privacy, we believe this line of research could assist in understanding diverse individuals’ mental health challenges and developing personalized interventions that improve the well-being and mental health of under-resourced communities Proferes et al. (2021).
Acknowledgment
We acknowledge the limited use of OpenAI’s ChatGPT in the writing and refining this manuscript.
@@ Appendix B Age and Gender
Figure 3: Differences in Covariates before and after CEM for groups “Control-ROW” and “MH-ROW”
Appendix C Coarsened Exact Matching
A control group is considered balanced with the treatment group if the difference is close to zero. The focus group MH-India has 1200 users and a total of 1200 users were matched from the groups MH-ROW and Control-ROW. Matching was not performed for Control-India as the number of samples is 930 and we did not want to drop any samples.
Appendix D Topic Uniqueness Scores
Table A3 describes the TU scores generated by each of the tested combinations of number of topics and grouping of Posts and Comments. This was used to choose Posts Only, Number of Topics = 2000 for our analysis.
Message Type Number of Topics Value
Both Messages 200 0.7205
Both Messages 500 0.6194
Both Messages 2000 0.5676
Posts Only 200 0.7675
Posts Only 500 0.7162
Posts Only 2000 0.7694
Table A3: Topic Uniqueness Scores for different numbers of topics and message types.
Appendix E Communication with Clinical Psychologists
@@
Table A4: Email communication with clinical psychologists who performed an informed review of the topics and ChatGPT-generated labels in this study.

### v3

@@ 4.1.2 LIWC
Fifty-two LIWC categories were significantly associated ($p<0.05$) with the MH-India group, whereas sixty categories were found to be correlated with the MH-RoW group. We provide the top LIWC categories for MH-India and MH-RoW in Fig 5 and 6. Sadness, a subset of negative emotions is more strongly correlated with MH-India whereas anxiety (also a type of negative emotion) is more correlated with MH-RoW. Mental health-related discourse is more present-focused in India whereas in MH-RoW group, it is past-focused. This aligns with n-grams findings i.e. MH-India’s emphasis on seeking help and MH-RoW’s on discussing symptoms. Social aspects including the use of 2nd/3rd person pronouns (e.g. you, we), communication, work, politeness, and achievement are positively correlated with only MH-India – reflecting the potential sociocultural expectations (such as collectivism and high-power distance in communication; Robert et al. (2000)) in Indian society.
Discussions in MH-RoW use more clinical language with concepts from psychosocial categories such as mental, health, and illness (also observed in Fig. A1) - reflecting awareness about mental health disorders. Substances are the most strongly correlated psychosocial category with MH-RoW however, it has a negative correlation with MH-India. The correlation with 1st person singular pronouns is also three times more in MH-RoW compared to MH-India. This aligns with recent findings that the use of self-referential pronouns is more likely a marker of depression in European Americans compared to other demographic groups. Similarly, cognitive processes such as all nor none (also a type of cognitive distortion Mercan et al. (2023)) and certitude are more strongly correlated with MH-RoW.
Categories such as illness and friend are similarly correlated with both groups.
[…]
@@ 4.1.3 LDA Topics
Of $2000$ topics, $109$ were found to be significant ($p<0.05$) for the MH-India group and $216$ for MH-RoW group. The top topics and the corresponding Pearson r, are provided in Fig. 7. Academic stressors (e.g. college, exam, study, university; learning, learn, data, programming) are exclusively related to discussions in MH-India. Family (family, mother, mom, father) and sexual content (porn, days, nofap, fap) are more frequent in MH-India compared to MH-RoW. Other topics including negative/suicidal thoughts (life, parents, family, hate, die), lack of belongingness (friends, talk, social, anxiety, alone; said, friend, told, friends), depression mentions (anymore, depression, tired, depressed; anxiety, depression, mental, medication) are overlapping however, they are more strongly correlated with MH-RoW. There are several topics (ants, colony, thingsforants, ant; blueface, snoop, dogg, gboard, etc,) which appear unrelated to mental health discussion but are discussed in both groups.
@@ 4.2 Validation
To what extent the open vocabulary topics having a significant correlation with the MH-India group are prevalent in Indian patients? - A Likert scale of 0-5 is provided where $5$ indicates ‘Highly Prevalent’ and ‘0’ indicates ‘Not observed at all’.
@@ 4.3 Is the language in MH-India different from control groups?
We used a logistic regression model with each feature set (i.e., n-grams, LIWC, and LDA topics) to predict membership in MH-India vs MH-RoW. High AUC scores (See Table 3) demonstrate that users’ language in the MH-India group significantly differs from those in the control groups, including MH-RoW. All language feature groups (i.e., n-grams, LIWC, and LDA topics) have fairly high AUC, indicating differences in mental health expressions at 1-3 gm level, in psycho-social categories as well as in latent topics of discussion.
@@ 5 Discussion
Our study uncovers significant cultural differences in the way users in the MH-India and MH-RoW groups discuss their mental health, underscoring the influence of sociocultural settings on how people perceive mental disorders and seek help. While MH-RoW discussions predominantly focus on feelings, symptoms, and peer support, MH-India situate their struggles with family, education, and work pressures. Although a similar distribution for age and gender is maintained across groups, academic and family pressures are unique to Indian users.
Another contrast observed in Fig 5 is the present-focused language used by Indian social media users, who often express immediate hopelessness or sadness with their current situation Bahri et al. (2023). This differs from the in-person Indian patients, who may ruminate more on past failures or anticipate future challenges. Pendse et al. (2019) also found that Indians discuss “wanting or needing friends” on Mental Health Support Forum more than other countries. We also observe similar behavior in MH-India on Reddit – Indian users are more likely to seek help on social media platforms (see Fig. 4 and 5). Similarly, somatization is the tendency to express emotional distress through bodily symptoms, such as fatigue, headaches, and body pain. This phenomenon is more prevalent in non-Western cultures, due to the heavy emphasis on addressing the tangible physical symptoms compared to the non-tangible mental ones Kirmayer and Young (1998). While this is widely observed among in-person patients in India, we observe a stronger correlation for health-related mentions (such as pain, sick, etc.) with MH-RoW (see Fig. 5 and 7). We speculate that this is likely due to the comorbidity of depression with physical disorders.
Finally, the solution-oriented communication style in MH-India, compared to the situation-descriptive posts in MH-RoW, raises critical questions about the predictive utility of social media language to monitor mental health risks in non-Western contexts. A similar observation is made for Facebook posts where Rai et al. (2024) noted that key markers of depression that is, the use of self-referential language and negative emotions on social media do not reflect depression in Black individuals while it does in White individuals.
@@
Regarding latent themes in Reddit discussions, only $56\%$ of 109 LDA topics correlated with the MH-India group were labeled as prevalent in Indian patients by clinical psychologists. We speculate that some of the topics not labeled as "prevalent" are unseen or emerging themes. Of Top-20 topics in MH-India, "not prevalent" topics revolve around Video Games/Hobbies and Programming/Learning indicating the influence of digital content and growing isolation amongst the undiagnosed young population. These topics could be underrecognized concerns. we also found topics related to Environmental Impact of Energy Sources, Humorous content and reactions, among others. Previous research has suggested that people (particularly young people) are increasingly climate anxious and that humor on social media is often used to cope with mental health challenges Schneider (2018); Sanson (2022). Furthermore, certain pop culture references may be crucial to understanding the narrative - as one clinical psychologist pointed out, the top word "Singh" in one of the topics may correspond to the suicide of late Bollywood star Sushant Singh Rajput, a significant event that potentially catalyzed a range of important conversations surrounding mental health in India Akbar et al. (2020).
@@
The growing treatment gap for mental disorders is a major concern in Indian society. The economic loss from mental health conditions between 2012-2030 is estimated at USD 1.03 trillion***United Nations: https://www.who.int/india/health-topics/mental-health. Automated systems that could diagnose and support mental well-being can potentially alleviate the lack of resources, but they would only be useful when designed considering the cultural sensitivities and norms of society. Based on our findings, culturally-aware mental health intervention techniques should seek to bridge the gap to treatment by addressing hypochondriacal ideas and familial embarrassment in particular. This study establishes significant linguistic variations in the mental health-related language in social media posts by Indians compared to individuals from the rest of the world. It thus becomes imperative to test AI models for cultural invariance before their deployment in clinical settings Aguirre et al. (2021); Rai et al. (2024).
@@ 6 Background
Depression and anxiety disorders are the most imminent mental health challenges, with the highest contribution to Indian Disability Adjusted Life Years  Sagar et al. (2020). Fear of shame is a primary barrier to mental health recovery in India whereas, for example, substance abuse is the major hurdle in America  Biswas et al. (2016). $71\%$ of Indians exhibited stigma when answering questions about mental health Foundation (2018). Relatedly, somatic symptoms, hypochondriasis, anxiety, and agitation are more commonly seen in Indian patients compared to psychological symptoms than British patients  Gada (1982). While the extent of mental health stigma and treatment (un)availability is often studied, it remains unknown how individuals suffering from mental disorders in India express and seek support on social media.
There is accumulating evidence that suggests language markers of depression vary with demographics such as race Rai et al. (2024); Aguirre and Dredze (2021), immigrant status Mittal et al. (2023a), and geographic location De Choudhury et al. (2017). For instance, Indian and South Africa-based users are less candid in their posts and less likely to exhibit negative emotions in comparison to their Western counterparts De Choudhury et al. (2017). Another study studied the language of Asian users including Indian, Malaysian, and Filipino on Mental Health Support Forum, Talklife Pendse et al. (2019) and found that Indians discuss “wanting or needing friends” more than other countries.
@@
While Reddit data is public, it may contain users’ personal information, including city and town. We limited our analysis to country and state-level geolocation information to reduce the possibility of personally identifying individuals. Gender was predicted using a continuous scale, with extremes indicating masculinity and feminity. We exercised caution while presenting linguistic patterns and examples not to reveal any individual’s timeline quotes. Members of our team have not viewed or worked with individual-level granular data. When done ethically with respect to user anonymity and privacy, we believe this line of research could assist in understanding diverse individuals’ mental health challenges and developing personalized interventions that improve the well-being and mental health of under-resourced communities.
@@ Appendix B Age and Gender
Figure A1: Top 25 1-3 grams in order of decreasing Pearson r for MH-India and MH-RoW.
Appendix C Communication with Clinical Psychologists
@@
Table A3: Email communication with clinical psychologists who performed an informed review of the topics.